OpenStudy (anonymous):

Find the y-coordinate of the vertex of the parabola whose equation is y = x^2 - x + 2. 2 1/4?

OpenStudy (anonymous):

Find the x-value of the vertex from x = -b/2a Then compute f(x) x = 1/2 y = 1/4 - 1/2 + 2 = ?

OpenStudy (anonymous):

Best advice: don't just look at a list of 4 choices and guess. Instead, compute!

OpenStudy (anonymous):

theres no choices in my textbook.. =2?

OpenStudy (anonymous):

Sorry, I just supposed that. Seems like most people have multiple choice. No choices is good in my opinion.

OpenStudy (anonymous):

oh i got the answer:) its 1 3/4 right?

OpenStudy (anonymous):

Yes.